# Stochastic gravitational-wave background search using data from five pulsar timing arrays

### 23 Mar 2026 v2

@@ Abstract
We deliberately kept this analysis straightforward and did not “iterate” it. This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised. Here, we avoid iterative refinement and tuning of the full PTA search pipeline. Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
@@
Notes: (1) Our full analysis code will be released publicly upon publication. (2) After this work was circulated, a different data-combination method was proposed [67]. Unlike ours, this does not use a physically consistent timing model for a given pulsar. See also SM Sec. I.
@@
Here, we provide additional plots and tables as supporting material.
I Effects discarding “outlier backend” TOAs
Table 4: Summary of EFAC cuts and their impact on backends and TOAs.
EFAC Backends Backends TOAs TOAs % TOAs
range discarded retained discarded retained discarded
$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
A recently posted paper [67] comments about “anomalous WN parameters” in a preliminary draft of this work. These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends. This was also the reason that 5 pulsars were given “special” treatment there. These phase disconnection issues were fixed by adding rotation counts to the input files.
II False-alarm probability $p$-value distributions
In the paper, Figure 3 shows a histogram of the posterior values of the false-alarm probability for the NPMV detection statistic. Here, in Figures 7 and 8, we provide the corresponding posterior $p$-value histograms for the Neyman-Pearson (NP) statistic and the traditional “optimal statistic” (OS). The means, medians, and widths of these distributions are reported in Table 3 of the paper.
III Data Analysis Workflow

### 28 Apr 2026 v3

@@ Abstract
Notes: (1) Our full analysis code will be released publicly upon publication. (2) After this work was circulated, a different data-combination method was proposed [60]. Unlike ours, that does not use a single consistent astrophysical/noise model for each pulsar. See also SM Sec. A.3.
Appendix A End Matter
EM1: The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the “uncorrelated” in CURN. It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms. Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR. Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2]. CURN is a more “conservative” null hypothesis than that of “no GWs”.
EM2: Some previous work determined $p$ values using “sky scrambling” and “phase shifting”, which were intended to “remove any signal from the data”. The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [61, 62]. The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [62]. However, neither of these methods properly reproduces the null distribution of $D$ [33]. For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
EM3: The data are merged as follows (also see SM Sec. A.1). If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties. The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
  1. 1.
  2. 2.
  3. 3.
    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [45] tools. See SM Sec. A.1.
  4. 4.
  5. 5.
EM4: The WN covariance matrix $\mathbf{N}$} characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs. 7.2 and 7.26], modeled with 30 and 100 frequency bins respectively. Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values. This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more precise estimates of the WN parameters. In SM Sec. A.3, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
EM5: NUTS is the default sampler for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers. It generates samples that are much more independent than the parallel tempering MCMC sampler [63] previously used by the individual PTAs, and is also much faster.
EM6: The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28]. The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28]. The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions, making it less sensitive to errors in modeling pulsar noise [28].
EM7: Here, the evidence $E$ is the likelihood marginalized over the prior [64]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec. A.3 Table 5). Evaluating the evidence integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$). To compute the evidences we use generalized stepping-stone sampling [65, 66, 67], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region. This gives $\ln(\rm{s}^{N}\,E_{\rm{HD}})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(\rm{s}^{N}\,E_{\rm{CURN}})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$.
EM8: Our HD reconstruction uses the NG [9] choice of 15 angular bins. The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs. For each posterior sample, the correlations of the pulsar pairs are weighted as in [57], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin. For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions. The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked. Fig. 3 shows their means (close to medians) and standard deviations.
@@ Appendix A End Matter
EM9: We deliberately kept this analysis straightforward and did not “iterate” it. This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised. Here, we avoid iterative refinement and tuning of the full PTA search pipeline. Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
@@
Here, we provide additional plots, figures, and tables as supporting material.
A.1 Data Analysis Workflow
@@
A.2 False-alarm probability $p$-value distributions
Here, we show histograms of the posterior values of the false-alarm probability for three different detection statistics.
We begin, in Figure 5, with the posterior $p$-value histograms for the NPMV statistic. The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$. Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances. The most likely range is shown in Fig. 5 as the band that contains 68% of the posterior points about the median. These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
Figures 6 and 7, are the corresponding histograms for the Neyman–Pearson (NP) statistic and the traditional “optimal statistic” (OS). The means, medians, and widths of these distributions are reported in Table 3 of the paper.
Why do we consider different statistics? Normally, the NP statistic would be preferred. However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise. In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them less sensitive to errors in modeling of the pulsar noise. Details may be found in [28].
A.3 Effects discarding “outlier backend” TOAs
Table 5: Summary of EFAC cuts and their impact on backends and TOAs.
EFAC Backends Backends TOAs TOAs % TOAs
range discarded retained discarded retained discarded
$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
A recently posted paper [60] comments about “anomalous WN parameters” in a preliminary draft of this work. These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends. This was also the reason that 5 pulsars were given “special” treatment there. We fixed these phase disconnection issues by by adding rotation counts to the input files (see A and SM Sec. A.1).